OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0 The solutions are -6,4 The equation is ___=0

OpenStudy (anonymous):

start with \[(x+6)(x-4)=0\] and then multiply out

OpenStudy (anonymous):

Is it x^2+6x+24 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[(x+6)(x-4)=x^2-4x+6x-24=x^2+2x-24\]

OpenStudy (anonymous):

Oh wow...thanks!

OpenStudy (anonymous):

So its x^2+2x-24?